When performing capital budgeting and considering replacement projects, one factor that must be considered is the potential __________ of equipment that is no longer needed.

a) taxation

  • b) salvage value
  • c) sunk costs
  • d) depreciation

Select one disadvantage of IRR as a capital budget method.

  • a) It is not useful for comparing projects with different lifespans.
  • b) It can only be used with projects that have positive cash flows.
  • c) It can be difficult to interpret and understand.
  • d) It fails to account for the time value of money.

With respect to payroll disbursements, one way a company can manage their cash more efficiently is to __________.

  • a) limit outsourcing
  • b) implement check kiting
  • c) increase float time
  • d) use lockbox banking

Which of the following is an example of a market risk for a company that manufactures automobiles?

  • a) A downgrade in the company's credit rating
  • b) Supply chain disruptions due to civil war in a country that supplies material
  • c) A massive lawsuit against the manufacturer over worker safety
  • d) A drop in demand due to the rise of ride-sharing as an alternative to automobile ownership

Consider the following data from a company's 95-day operating cycle:

  • Payable days: 8
  • Receivable days: 25
  • Inventory days: 70

What is the cash conversion cycle for this company?

  • a) 87
  • b) 103
  • c) 47
  • d) 53

Farrah owns 500 shares of stock valued at $30/share in Company A.

After the company issues a 3% stock dividend, what does Farrah own?

  • a) 500 shares valued at $30.90/share
  • b) 515 shares valued at $30/share
  • c) 500 shares valued at $30/share
  • d) 515 shares valued at $29.13/share
Answer #1

1) The correct answer is

b) Salvage value

Salvage value should be considered because this will reduce the net initial capital outflow.

2) The correct answer is (a)

It is not useful for comparing projects with different lifespans

Other options are not correct. For IRR there has to be at least change of sign once. IRR does consider time value of money.

3) The correct answer is (d)

Use Lockbox banking

Lockbox banking can reduce the processing time significantly.

4) The correct answer is (b)

Supply chain disruptions due to civil war in a country that supplies material

Other options are risk related one specific industry or company, so they would be non -systematic risk.

5) Cash conversion cycle = Inventory days + receivable days – payable days

70 + 25 – 8 = 87

6) the stock dividend does not increase the overall value

So, number of shares = 500*1.03 = 515

Price would decrease to 30/(1+0.03) = 29.13